OpenStudy (anonymous):

What is the solution set for |–x + 2| = 5?

OpenStudy (ajprincess):

|–x + 2| = 5 => -x+2=5 or -x+2=-5 Solve for x in both cases. Does that help? @Ashleytaylor2014

OpenStudy (anonymous):

yes thank you!

OpenStudy (ajprincess):

Welcome:)